# On an approach to canonicalizing elliptic Feynman integrals

### v3

@@ Appendix D Special relations of elliptic integrals
In this appendix, we will briefly introduce some special relations that changes the arguments of elliptic integrals. We have changed the arguments $\lambda=4k/(1+k)^{2}$ to $k^{2}$ from Eqs. (95) to Eqs. (96), and we know the two kinds of arguments are the ones for Legendre family in Sec. 2 and Jacobi family in App. B, respectively. A natural way to relate the elliptic integrals with these two kinds of arguments is to consider the Möbius transformation Eq. (147) which bridges them, combined with the quadratic transformation in Eq. (17).
@@
We can also use the periods defined in Eqs. (26) and Eqs. (129) to rewrite the relations above, which are
@@
These arguments correspond to the six cross ratios for a single elliptic curve. Moreover, since the LHS of the relations are fixed, such as $K(k^{2})$ and $K(1-k^{2})$, we can also derive relations for elliptic integrals whose arguments are among those in (203). These relations can be interpreted as reflecting different choices of the modulus $\lambda$ for a single elliptic curve. This interpretation can be verified by examining the Klein’s j-invariants.
In Eqs. (199), there are two more different arguments $k^{2}$ and $1-k^{2}$ on the LHS which do not belong to (203), implying that these relations connect two distinct elliptic curves. However, there exist specific relationships between these two elliptic curves. As shown in App. B.2.3, the two elliptic curves are connected via a quadratic transformation, up to a Möbius transformation, which does not alter the elliptic curve. Consequently, the two elliptic curves are isogenic, as evidenced by
@@
where we’ve applied Eqs. (199) with arguments in Eqs. (159) and (93) for a concrete example.

### v4

@@ Appendix D Special relations of elliptic integrals
In this appendix, we will briefly introduce some special relations that changes the arguments of elliptic integrals. We have changed the arguments $\lambda=4k/(1+k)^{2}$ to $k^{2}$ from Eq. (95) to Eq. (96), and we know the two kinds of arguments are the ones for Legendre family in Sec. 2 and Jacobi family in App. B, respectively. A natural way to relate the elliptic integrals with these two kinds of arguments is to consider the Möbius transformation Eq. (147) which bridges them, combined with the quadratic transformation in Eq. (17).
@@
We can also use the periods defined in Eqs. (26) and (129) to rewrite the relations above, which are
@@
These arguments correspond to the six cross ratios for a single elliptic curve. Moreover, since the LHS of the relations are fixed, such as $K(k^{2})$ and $K(1-k^{2})$, we can also derive relations for elliptic integrals whose arguments are among those in Eq. (203). These relations can be interpreted as reflecting different choices of the modulus $\lambda$ for a single elliptic curve. This interpretation can be verified by examining the Klein’s j-invariants.
In Eq. (199), there are two more different arguments $k^{2}$ and $1-k^{2}$ on the LHS which do not belong to Eq. (203), implying that these relations connect two distinct elliptic curves. However, there exist specific relationships between these two elliptic curves. As shown in App. B.2.3, the two elliptic curves are connected via a quadratic transformation, up to a Möbius transformation, which does not alter the elliptic curve. Consequently, the two elliptic curves are isogenic, as evidenced by
@@
where we’ve applied Eq. (199) with arguments in Eqs. (159) and (93) for a concrete example.